OpenStudy (majacw):

write an equation for the line that is parallel to the given line and passes through the given point. y=4x+7;(3,11)

OpenStudy (anonymous):

So start with what you know about parallel lines. (hint: slope)

OpenStudy (anonymous):

Parallel lines have the same slopes.... right?

OpenStudy (majacw):

yes

OpenStudy (anonymous):

Ok.... so anyway you take the point and plug it into an equation (we're now finding the y intercept): 11=4(3)+b

OpenStudy (anonymous):

solve that please.... tell me what you get

OpenStudy (majacw):

u dont have to give me the answer, just help to get there

OpenStudy (anonymous):

Ok... well in order to get there you gotta solve for b with the equation above.... then you use your slope and get y=4x+(what you get for b)

OpenStudy (anonymous):

so tell me what you get.... and I'll tell you if it's right

OpenStudy (anonymous):

did you get your answer?

OpenStudy (anonymous):

Well... I gotta go but if you wanna make sure that you're right the answer is y=4x-1

OpenStudy (majacw):

im sorry, mom called can you go on

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Do you understand how I got that?

OpenStudy (majacw):

I do not, I'm sorry.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Let me explain what I did all at once

OpenStudy (anonymous):

So parallel lines have the same slope... so your slope is 4x.... what you need to do to make it the line on point (3,11) is find the y intercept... so you set up this equation: 11=4(2)+b -12 -12 -1=b so now put it all together: y=4x-1

OpenStudy (anonymous):

does that make more sense now?

OpenStudy (majacw):

yes and no. i think that I just need to study more and get some one on one help. thank u though.

OpenStudy (anonymous):

Wait what doesn't make sense?... I'm pretty good at explaining... maybe I can help some more?

OpenStudy (majacw):

lol, it's not you or how you explained, its going over my head a little bit. this is all kinda hard to begin with for me. i have a tutor but wont see him for a few days.

OpenStudy (anonymous):

oh ok! good luck! I'm sure you'll get the hang of it pretty fast after you catch up with yout tutor!

OpenStudy (anonymous):

I mean your.... lol

OpenStudy (anonymous):

Bye!

OpenStudy (majacw):

thank you and sorry if i wasted your time

OpenStudy (anonymous):

No problem! :-)